Please help me with this algebra problem. Thanks so much in advance.
Simplify and show all work please
12 divided by 1/4 + 2/5 - 2^3

Answer by rfer(16322)

12*4/1+2/5-8=
48+2/5-8=
40+2/5=40 2/5
200/5+2/5=202/5=40 2/5